OpenStudy (anonymous):

I will medal/fan.. NO DIRECT ANSWERS Identify the type of sequence shown in the table below and select the appropriate response. n f(n) 1 540 2 500 3 460 4 420 5 380 Arithmetic sequence; common difference is −40 Arithmetic sequence; common difference is 40 Geometric sequence; common ratio is −40 Geometric sequence; common ratio is 40

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

Do you know what each of those sequences look like?

OpenStudy (anonymous):

somewhat?! @DanJS

OpenStudy (anonymous):

yes.. okay! @DanJS @Hero

OpenStudy (danjs):

Test if it follows each by using a couple of the points

OpenStudy (anonymous):

I am ust not good with the arithmatic stuff like what?

OpenStudy (danjs):

If it increases the same amount for each increase in n, it is arithmetic

OpenStudy (danjs):

or decreases

OpenStudy (anonymous):

so Its geometric, I am assuming?

OpenStudy (danjs):

a geometric sequence will have the ratio of any two consecutive terms as a constant same value

OpenStudy (anonymous):

so the answer is c or d, not a or b.

OpenStudy (danjs):

looks like it is decreasing by 40 for every increase in n

OpenStudy (danjs):

arithmetic, common difference in sucessive terms

OpenStudy (anonymous):

okay. so I mean a or b.. I was reading it wrong

OpenStudy (anonymous):

so the common difference is -40?

OpenStudy (danjs):

yes, you see you are taking the previous term and adding (-40) to it every time

OpenStudy (danjs):

an = a(n-1) - 40

OpenStudy (anonymous):

Can you help with another?

OpenStudy (danjs):

If it were geometric, you can test out the ratio to see, by \[\frac{ a _{n+1} }{ a _{n} } = r\]

OpenStudy (danjs):

the ratio will be a value for geometric, not the difference

OpenStudy (anonymous):

What is the exact value for the expression the square root of 56. − the square root of 14. + the square root of 126.? Simplify if possible. the square root of 14. 4 the square root of 14. 2 the square root of 42. 8 the square root of 42.

OpenStudy (anonymous):

sorry dinners ready... Maybe later, if ur on you could help? ::)

OpenStudy (anonymous):

@DanJS I think that the answer is B?